The Earth's atmosphere at sea level and under normal conditions has a pressure of 1.01x10^5 Pa, which is due to the weight of the air above the ground pushing down on it. How much force due to this pressure is exerted on the roof of a building whose dimensions are 10.2 m long and 14.5 m wide?
Answer #1

Force on an area A due to a pressure P is given by

F=PA

Area of the roof is

A=\left (10.2\times14.5 \right )\ m^2=147.9\ m^2

hence, force on the roof of the building is

F=1.01\times10^5\times147.9\ N=14937900\ N